40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5465     Entry time: Mon Sep 19 16:56:29 2011
Author: Paul 
Type: Update 
Category: SUS 
Subject: SRM oplev pitch calibration 

 Same measurements for SRM pitch (as previously done for yaw in entry 5460) are complete. The QPD is back in the path and aligned. I will be doing the same measurements for ITMY now though, so please ask before activating the SRM or ITMY oplev servos, as I may be blocking the beam.

ELOG V3.1.3-